OpenStudy (anonymous):

what is the sum of the roots of the polynomial? x^3+2x^2-11x-12

OpenStudy (yttrium):

where are you stocked?

OpenStudy (anonymous):

well I used synthetic division to solve this problem like someone else suggested, but when I did the answer was wrong so I'm confused on how to solve it, or maybe there was a error in my process... but other then that I don't really know.

hartnn (hartnn):

sum of roots of a cubic equation, \(\Large ax^3+bx^2+cx+d= 0\) is just -b/a

OpenStudy (yttrium):

Oh really? @hartnn How about if there's no x^2 term? Is it zero?

hartnn (hartnn):

then b=0 sum = 0

OpenStudy (yttrium):

Ohh thanks! :)) How about the sum of roots of other degrees? Are there any shortcut?

hartnn (hartnn):

yes

hartnn (hartnn):

\(\Large ax^n +bx^{n-1}+..... = 0\) sum of roots always = -b/a

OpenStudy (yttrium):

Ohhh awesome! :) Thanks

OpenStudy (whpalmer4):

Cool! That's today's installment of "learn something new every day on OpenStudy"...

hartnn (hartnn):

\(Roots : p,q,r \\ ax^3+bx^2+cx+d=0 \\ p+q+r = -b/a \\ pq+qr+pr = c/a\\ pqr = - d/a\)

OpenStudy (anonymous):

lol thanks to everyone :)

OpenStudy (whpalmer4):

\[pqr = -d/a\]is obvious enough, but the others might take a bit of rumination...

hartnn (hartnn):

\(Roots : p,q,r,s \\ ax^4+bx^3+cx^2+dx+e=0 \\ p+q+r+s = -b/a \\ (roots ~taken ~2~at~a~time) = c/a\\ (roots ~taken ~3~at~a~time) = - d/a \\ (roots ~taken ~4~at~a~time) = pqrs = e/a \)

hartnn (hartnn):

in general, such pattern is true for any value of n

OpenStudy (yttrium):

what does roots taken 2 at a time means??

hartnn (hartnn):

pq+pr+ps+qr+qs+rs

OpenStudy (yttrium):

Ohh okay.